OpenStudy (anonymous):

-27<1+7p suppose to be a line under this <

jimthompson5910 (jim_thompson5910):

-27 <= 1+7p -27-1 <= 7p -28 <= 7p -28/7 <= p -4 <= p p >= -4 So the answer is p is greater than or equal to -4

OpenStudy (anonymous):

which way is th arrow suppose to be going

jimthompson5910 (jim_thompson5910):

it's going to the right since the sign is greater than or equal to

OpenStudy (anonymous):

is it suppose to be going right

OpenStudy (anonymous):

k

OpenStudy (anonymous):

can you help me with other one

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):